# Framework for Variable-lag Motif Following Relation Inference In Time Series using Matrix Profile analysis

### 05 Jan 2024 v1

@@ Abstract.
Knowing who follows whom and what patterns they are following are crucial steps to understand collective behaviors (e.g. a group of human, a school of fish, or a stock market). Time series is one of resources that can be used to get insight regarding following relations. However, the concept of following patterns or motifs and the solution to find them in time series are not obvious. In this work, we formalize a concept of following motifs between two time series and present a framework to infer following patterns between two time series. The framework utilizes one of efficient and scalable methods to retrieve motifs from time series called the Matrix Profile Method. We compare our proposed framework with several baselines. The framework performs better than baselines in the simulation datasets. In the dataset of sound recording, the framework is able to retrieve the following motifs within a pair of time series that two singers sing following each other. In the cryptocurrency dataset, the framework is capable of capturing the following motifs within a pair of time series from two digital currencies, which implies that the values of one currency follow the values of another currency patterns. Our framework can be utilized in any field of time series to get insight regarding following patterns between time series.
@@ 2. Related work
The motif discovery problem was introduced in the work (Chiu et al., 2003), and since then there are several methods have been proposed to find motifs in time series (Van Onsem et al., 2023; Zhu et al., 2016; Liu et al., 2023; Li and Lin, 2010; Linardi et al., 2018). One of the well-known efficient methods to retrieve motifs from time series is a Matrix Profile framework introducing in (Yeh et al., 2016; Zhu et al., 2016). The matrix profile framework was developed to discover motifs in time series in (Zhu et al., 2016) called ”STOMP”. The framework is scalable to compute million data points to find exact motifs in reasonable time, which makes in one of the efficient framework to deploy for finding motifs. While matrix profile can be efficiently and scalably utilized to find many types of patterns in time series (e.g. motifs (Zhu et al., 2016), discord (Yeh et al., 2016), Shapelets (Yeh et al., 2016), anomaly patterns (Tafazoli and Keogh, 2023; Lu et al., 2022)), to the best of our knowledge, the matrix profile has not been utilized to solve the problem of finding following motifs yet.
In this work, we formalize Variable-Lag Following Motifs Inference Problem and propose a framework to find a following motifs in time series based on the matrix profile method in (Zhu et al., 2016). In the next section, the problem formalization and related details are provided.
@@ 4. Methods
4.1. Matrix Profile’s algorithm
@@
To find the most similar parts of two inputted time series, Matrix Profile algorithm results in Matrix Profile (MP), which represents z-normalized euclidean distance between the most similar two parts of the time series, and Matrix Profile Index (MPI), which represents time steps of the most nearest part of the inputs.
@@ 9 $MPI[i]\leftarrow i$;
Algorithm 3 Matrix Profile’s Algorithm
Matrix profile algorithm does window sliding with reference time series by a window-size query, which cuts from the query time series, to calculate z-normalized euclidean distance in each slide. Thus, one sliding of the query results in a distance value for each time step of the reference time series. Excluding trivial matching time steps, Matrix profile and Matrix profile indices array are updated the lower distance value if the distance value is lower than their current element. After the matrix profile algorithm finishes every sliding window of cutting query and calculating distance for each query and reference time series. Then, Matrix profile and the indices array, which has its size as the length of reference time series subtracted with window size, contain the shortest distances and the most similar query starting indices.
@@ 12 Return $result_{lead},result_{leadvalue},index_{motif}$;
In the aspect of computational cost, the proposed algorithm has time complexity as $O(n^{2}log(n))$ since we use STAMP as the main computation part where $n$ is the length of time series.
@@ 7. Conclusion
In this work, we formalized a concept of following motifs between two time series and presented a framework to infer following patterns from data. The framework utilized one of efficient and scalable methods to retrieve motifs from time series called matrix profile method. We compared our proposed framework with several baselines. The framework performed better than baselines in the simulation datasets. In the dataset of sounds recording, the framework was able to retrieve the following motifs within a pair of time series that two singers sang following each other. In the cryptocurrency dataset, the framework was capable of capturing the following motifs within a pair of time series from two digital currencies, which implied that the values of one currency followed the values of another currency patterns. Our framework can be utilized in any field of time series to get insight regarding following patterns between time series. The code and datasets can be found at https://github.com/hughnaaek/Following-Motif-Relation.

### 18 Jan 2024 v2

@@ Abstract.
Knowing who follows whom and what patterns they are following are crucial steps to understand collective behaviors (e.g. a group of human, a school of fish, or a stock market). Time series is one of resource that can be used to get insight regarding following relations. However, the concept of following patterns or motifs and the solution to find them in time series is not obvious. In this work, we formalize a concept of following motifs between two time series and present a framework to infer following patterns from data. The framework utilizes one of efficient and scalable methods to retrieve motifs from time series called matrix profile method. We compare our proposed framework with several baseline. The framework performs better than baseline in the simulation dataset. In the dataset of sounds from video, the framework is able to retrieve the following motifs within a pair to time series that two singers sink following each other. In the cryptocurrency dataset, the framework is capable of capturing the following motifs within a pair of time series from two digital currency. Our framework can be utilized in any field of time series.
@@ 2. Related work
The motif discovery problem was introduced in the work (Chiu et al., 2003), and since then there are several methods have been proposed to find motifs in time series (Van Onsem et al., 2023; Zhu et al., 2016; Liu et al., 2023; Li and Lin, 2010; Linardi et al., 2018). One of the well-known efficient methods to retrieve motifs from time series is to use the representation called Matrix Profile introducing in (Yeh et al., 2016; Zhu et al., 2016). The matrix profile inference frameworks were developed to discover motifs in time series in (Yeh et al., 2016; Zhu et al., 2016). The frameworks were scalable to compute million data points to find exact motifs in reasonable time, which makes in one of the efficient framework to deploy for finding motifs. While matrix profile can be efficiently and scalably utilized to find many types of patterns in time series (e.g. motifs (Zhu et al., 2016), discord (Yeh et al., 2016), Shapelets (Yeh et al., 2016), anomaly patterns (Tafazoli and Keogh, 2023; Lu et al., 2022)), to the best of our knowledge, the matrix profile has not been utilized to solve the problem of finding following motifs yet.
In this work, we formalize Variable-Lag Following Motifs Inference Problem and propose a framework to find a following motifs in time series based on the matrix profile method in (Yeh et al., 2016). In the next section, the problem formalization and related details are provided.
@@ 4. Methods
4.1. Matrix Profile inference
@@
To find the most similar parts of two inputted time series, Matrix Profile inference algorithm results in Matrix Profile (MP), which represents z-normalized Euclidean distance every subsequence and its (non-trivial) nearest neighbor within the same time series, and Matrix Profile Index (MPI), which represents time steps of the most nearest part of the inputs.
@@ 9 $MPI[i]\leftarrow i$;
Algorithm 3 Matrix Profile inference algorithm
To illustrate how to perform matrix profile inference, the Matrix profile inference algorithm 3 is one example (STAMP) of many algorithms that can perform the task. Algorithm 3 does window sliding with reference time series by a window-size query, which cuts from the query time series, to calculate z-normalized euclidean distance in each slide. Thus, one sliding of the query results in a distance value for each time step of the reference time series. Excluding trivial matching time steps, Matrix profile and Matrix profile indices array are updated the lower distance value if the distance value is lower than their current element. After the matrix profile algorithm finishes every sliding window of cutting query and calculating distance for each query and reference time series. Then, Matrix profile and the indices array, which has its size as the length of reference time series subtracted with window size, contain the shortest distances and the most similar query starting indices.
@@ 12 Return $result_{lead},result_{leadvalue},index_{motif}$;
In the aspect of computational cost, the proposed algorithm has time complexity as $O(n^{2}log(n))$ since we use STAMP (Yeh et al., 2016) as the main computation part where $n$ is the length of time series. Moreover, note that we could also use SCRIMP++ (Zhu et al., 2018) which is only $O(n^{2})$. Both STAMP and SCRIMP++ allow computation both as anytime and contract algorithms (López-Ortiz et al., 2014). In practice this implies that datasets with up to say ten million datapoints can easily be analyzed in realtime interactive sessions.
@@ 7. Conclusion
In this work, we formalized a concept of following motifs between two time series and presented a framework to infer following patterns from data. The framework utilized one of efficient and scalable methods to retrieve motifs from time series called matrix profile method. We compared our proposed framework with several baselines. The framework performed better than baselines in the simulation datasets. In the dataset of sounds recording, the framework was able to retrieve the following motifs within a pair of time series that two singers sang following each other. In the cryptocurrency dataset, the framework was capable of capturing the following motifs within a pair of time series from two digital currencies, which implied that the values of one currency followed the values of another currency patterns. Our framework can be utilized in any field of time series to get insight regarding following patterns between time series.
The code and datasets can be found at https://github.com/hughnaaek/Following-Motif-Relation.